# Supersaturation of Tree Posets

### v1

@@ Abstract
A well-known result of Bukh [6] shows that if $P$ is a tree poset of height $k$, then any subset of the Boolean lattice $\mathcal{F}\subseteq\mathcal{B}_{n}$ of size at least $(k-1+\varepsilon){n\choose\lceil n/2\rceil}$ contains at least one copy of $P$. This was extended by Boehnlein and Jiang [5] to induced copies. We strengthen both results by showing that for any integer $q\geq k$, any family $\mathcal{F}$ of size at least $(q-1+\varepsilon){n\choose\lceil n/2\rceil}$ contains on the order of as many induced copies of $P$ as is contained in the $q$ middle layers of the Boolean lattice. This answers a conjecture of Gerbner, Nagy, Patkós, and Vizer [13] in a strong form for tree posets.
@@ 1.1 History and Background
The celebrated Sperner’s theorem [25] in extremal set theory determines the size of the largest family of sets in $[n]$ not containing a $2$-chain $F_{1}\supset F_{2}$ Later, Erdős [12] extended Sperner’s theorem to determine the largest family not containing a $k$-chain $F_{1}\supset F_{2}\supset\cdots\supset F_{k}$. Katona and Tarján [18] then initiated a systematic study of the size of the largest set family in the Boolean lattice $\mathcal{B}_{n}$ (i.e. the poset on $2^{[n]}$ under the set inclusion relation) that avoids a given subposet. This topic has attracted much attention and witnessed many advances in the last decades, particularly in the last decade or so. We refer the interested reader to the nice survey paper by Griggs and Li [15].
We now formalize the problem precisely. Given two posets $P$ and $Q$, a poset homomorphism is a function $f:P\to Q$ such that $f(x)\leq f(y)$ whenever $x\leq y$. An induced poset homomorphism is a function $f:P\to Q$ such that $f(x)\leq f(y)$ if and only if $x\leq y$. We say that a poset $Q$ contains another poset $P$ if there is an injective poset homorphism from $P$ to $Q$. We say that a poset $Q$ contains an induced copy of another poset $P$ if there is an injective induced poset homomorphism from $P$ to $Q$. If a poset $Q$ does not contain a copy of another poset $P$, we say that $Q$ is $P$-free. If $Q$ does not contain an induced copy of $P$, we say that $Q$ is induced $P$-free. We define the Hasse diagram of $P$, denoted $H(P)$, as a directed graph with vertex set $P$ where there is an edge from $x$ to $y$ upwards only if $y>x$ and there is no $z$ such that $y>z>x$. We recall the following definitions for the largest size of a family which is (induced) $P$-free.
Definition 1.1.
Given a poset $P$ and an integer $n$, we define $La(n,P)$ to be the largest size of a $P$-free subfamily of $\mathcal{B}_{n}$ and $La^{*}(n,P)$ the largest size of an induced $P$-free subfamily of $\mathcal{B}_{n}$.
Using this notation, Sperner’s theorem [25] says that $La(n,C_{2})=La^{*}(n,C_{2})=\binom{n}{\lceil n/2\rceil}$ where $C_{2}$ is the chain of length 2. Motivated by a number of early results in the study of $La(n,P)$ (see for example [8, 9, 14, 27]), Griggs and Lu [16] and independently Bukh [6] made the following central conjecture.
@@
A similar conjecture for the induced $P$-free problem was formally stated in Gerbner, Nagy, Patkós, and Vizer [13].
Conjecture 1.3 (Gerbner-Nagy-Patkós-Vizer [13]).
Let $P$ be a poset. Then
$La^{*}(n,P)=(1+o(1))e^{*}(P)\binom{n}{\lfloor n/2\rfloor},$
where $e^{*}(P)$ denotes the largest integer $\ell$ such that for all $j$ and $n$ the family $\bigcup_{i=1}^{\ell}\binom{[n]}{i+j}$ is induced $P$-free.
These conjectures remained open until recently when they were proven to be false for $P$ the $d$-dimensional Boolean lattice with $d\geq 4$ due to a beautiful construction of Ellis, Ivan, and Leader [11]. Nevertheless, one can ask to what extent weaker versions of these conjectures hold. For example, one of the most general results regarding estimates for $La(n,P)$ and $La^{*}(n,P)$ is due to Methuku and Pálvölgyi [19], who showed that for every poset $P$, there exists a constant $C_{P}$ such that $La^{*}(n,P)\leq C_{P}\binom{n}{\lfloor n/2\rfloor}$ with the value of $C_{P}$ later being improved upon by Tomon [28].
In another direction, one can try and establish which posets $P$ satisfy the conjectures above. There are a number of posets which do this (see [15] for further details), with one particularly broad class of examples being tree posets, which are posets whose Hasse diagram is a tree. Indeed, a general result of Bukh [6] established Conjecture 1.2 for tree posets, with his result later being extended by Boehnlein and Jiang [5] to the induced setting of Conjecture 1.3.
@@
Theorem 1.5 (Boehnlein-Jiang [5]).
Let $P$ be a tree poset of height $k$. Then
$La^{*}(n,P)=\left(k-1+O\left(\frac{\sqrt{n\log n}}{n}\right)\right)\binom{n}{\lfloor n/2\rfloor}.$
1.2 Main Results
The main result of this paper is a general result in the form of Theorem 5.9 which says that any subset $\mathcal{F}\subseteq\mathcal{B}_{n}$ of size at least $(q-1+\varepsilon){n\choose\lfloor n/2\rfloor}$ contains a large “robust” sequence of $q$-chains. This result can be used to resolve a number of extremal problems for tree posets, but for brevity we discuss only one such application related to a conjecture of Gerbner, Nagy, Patkós, and Vizer [13] regarding supersaturation of posets.
As is typical in extremal combinatorics, once we develop some knowledge about the threshold for the occurrence of a certain substructure, it is natural to ask how many copies of the substructures we can guarantee once we go beyond this threshold; such problems are typically referred to as supersaturation problems. An early example of this kind was a question considered by Erdős and Katona, who conjectured that a family with $\binom{n}{\lfloor n/2\rfloor}+t$ sets in $[n]$ must contain at least $t\cdot\lfloor\frac{n+1}{2}\rfloor$ many $2$-chains. This conjecture was confirmed by Kleitman [17], who in fact showed that for every $0\leq a\leq 2^{n}$, every set family in $[n]$ with size $a$ contains at least as many $2$-chains as the so-called centralized family of size $a$, i.e. a family of $a$ sets whose cardinalities are as close to $n/2+1/4$ as possible. Kleitman [17] further conjectured that the same families should also minimize the number of $k$-chains for every $k$. Five decades later, Kleitman’s result was rediscovered by Das, Gan, and Sudakov [7] and independently by Dove, Griggs, Kang, and Sereni [10]. Both papers further confirmed Kleitman’s conjecture for every $k$ and $a$ belonging to a certain range above the sum of the $k-1$ largest binomial coefficients. Subsequently, Balogh and Wagner [4] proved Kleitman’s conjecture for all $k$ and $a\leq(1-\varepsilon)2^{n}$, provided that $n$ is sufficiently large with respect to $k$ and $\varepsilon$. Finally, Samotij [23] resolved Kleitman’s conjecture in full.
In the spirit of these works, it is natural to ask in general how many copies of a poset $P$ we can ensure in a set family in $[n]$ with size sufficiently larger than $La(n,P)$ or how many induced copies of $P$ we can ensure in a set family in $[n]$ with size sufficiently larger than $La^{*}(n,P)$. Besides the work mentioned above, Patkós [21] obtained some tight supersaturation results for the butterfly poset. Gerbner, Nagy, Patkós, and Vizer [13] formulated a general conjecture for this problem, and for this we introduce some definitions.
Definition 1.6.
Given a poset $P$ and integers $q,n$, let $M(n,q,P)$ denote the number of copies of $P$ in the $q$ middle levels of $2^{[n]}$ and let $M^{*}(n,q,P)$ denote the number of induced copies of $P$ in the $q$ middle levels of $2^{[n]}$.
Conjecture 1.7 (Gerbner-Nagy-Patkós-Vizier [13], Conjecture 4).
  1. 1.
    For every poset $P$ and $\varepsilon>0$ there exists $\delta>0$ such that if $\mathcal{F}\subseteq 2^{[n]}$ is of size at least $(e(P)+\varepsilon)\binom{n}{\lfloor n/2\rfloor}$, then $\mathcal{F}$ contains at least $\delta\cdot M(n,e(P)+1,P)$ many copies of $P$.
  2. 2.
    For every poset $P$ and $\varepsilon>0$ there exists $\delta>0$ such that if $\mathcal{F}\subseteq 2^{[n]}$ is of size at least $(e^{*}(P)+\varepsilon)\binom{n}{\lfloor n/2\rfloor}$, then $\mathcal{F}$ contains at least $\delta\cdot M^{*}(n,e^{*}(P)+1,P)$ many induced copies of $P$.
The non-induced part of Conjecture 1.7 was verified in [13] for tree posets which either have height 2 or which are monotone. However, since Conjecture 1.7 is a strengthening of Conjecture 1.2 and Conjecture 1.3, the previously mentioned work of Ellis, Ivan, and Leader [11], continues to show that Conjecture 1.7 is not true for $d$-dimensional Boolean lattice with $d\geq 4$. Still, it is an interesting question as to which families of posets this conjecture holds for. As the main result of our paper, we show that Conjecture 1.7 holds for tree posets $P$ in the following strong form.
@@ 1.2 Main Results
Let $k$ be a fixed positive integer. Let $P$ be a tree poset of height $k$. Then for any real $\varepsilon>0$ and integer $q\geq k$ there exists a real $\delta>0$ such that every family $\mathcal{F}\subseteq\mathcal{B}_{n}$ with $|\mathcal{F}|\geq(q-1+\varepsilon)\binom{n}{\lfloor n/2\rfloor}$ contains at least $\delta\cdot M(n,q,P)$ induced copies of $P$.
As a corollary to Theorem 1.8, we get the following explicit result which is not hard to check is tight for saturated tree posets, i.e. those for which every maximal chain has the same length.
@@
Let $k$ be a fixed positive integer. Let $P$ be a tree poset of height $k$. For any real $\varepsilon>0$, there exists a real $\delta>0$ such that every subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ with $|\mathcal{F}|\geq(k-1+\varepsilon)\binom{n}{\lfloor n/2\rfloor}$ contains at least $\delta n^{|P|-1}\binom{n}{\lfloor n/2\rfloor}$ induced copies of $P$.
2 Sketch of Proof
Due to the nature of the containment relation of $\mathcal{B}_{n}$, it is most convenient to associate each member $F$ of a given subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ with the full chains $\chi$ of $\mathcal{B}_{n}$ that contain it. So, we will be working with pairs $(\chi,F)$ instead of individual members $F$. Likewise, we will associate each $q$-chain $Q$ in our family $\mathcal{F}$ with the full chains $\chi$ of $\mathcal{B}_{n}$ that contain it and call such pairs $(\chi,Q)$ $q$-marked chains. This approach of working with $q$-marked chains was originated by Bukh [6] which was later further developed by Boehnlein and Jiang [5] to address induced copies of tree posets. Here, we push these ideas even further to ensure that we can find copies of a tree poset in a robust way.
With the specific goal of Theorem 1.8 in mind, we start by noting that there is no explicit formula for the number of copies of $P$ in the $q$ middle layers $M(n,q,P)$.
However, this is not a major problem since there are only a bounded number of ways of embedding $P$ into the middle $q$ levels of $\mathcal{B}_{n}$ in terms of which level each member of $P$ is mapped to. So it suffices if we can show that for any such level assignment $I:P\rightarrow[q]$ of members of $P$ to the middle $q$ levels, the number of copies of $P$ guaranteed in $\mathcal{F}$ is as least as large as a fraction of number of copies of $P$ to the middle $q$ levels obeying that particular level assignment.
At a high level, our strategy for embedding of $P$ is as follows. We fix an ordering $x_{1},x_{2},\ldots,x_{|P|}$ of vertices of $P$ such that each $x_{j+1}$ has a unique neighbour among $x_{1},\dots x_{j}$, which we call $x_{j}$’s parent $p(x_{j})$, and we fix an assignment $I$ as above. We embed the vertices $x_{1},x_{2},\dots$ iteratively, ensuring that at each step there are many choices to embed the current vertex $x_{j}$. To ensure this, we generate a nested family of $q$-marked chains $\mathcal{M}^{0}\supseteq\mathcal{M}^{1}\supseteq\cdots\supseteq\mathcal{M}^{|P|}$ such that $x_{j}$ is embedded in $I(x_{j})$th member of some $q$-marked chain in $\mathcal{M}^{|P|-j+1}$ and each $\mathcal{M}^{i+1}$ is “robust” with respect to $\mathcal{M}^{i}$ in the sense that given an embedding of $p(x_{j})$, there are “many” choices to embed $x_{j}$.

### v2

# Tree Posets: Supersaturation, Enumeration, and Randomness

@@ Abstract
We develop a powerful tool for embedding any tree poset $P$ of height $k$ in the Boolean lattice which allows us to solve several open problems in the area. We show that:
  • If $\mathcal{F}$ is a family in $\mathcal{B}_{n}$ with $|\mathcal{F}|\geq(q-1+\varepsilon){n\choose\lfloor n/2\rfloor}$ for some $q\geq k$, then $\mathcal{F}$ contains on the order of as many induced copies of $P$ as is contained in the $q$ middle layers of the Boolean lattice. This generalizes results of Bukh [9] and Boehnlein and Jiang [8] which guaranteed a single such copy in non-induced and induced settings respectively.
  • The number of induced $P$-free families of $\mathcal{B}_{n}$ is $2^{(k-1+o(1)){n\choose\lfloor n/2\rfloor}}$, strengthening recent independent work of Balogh, Garcia, Wigal [1] who obtained the same bounds in the non-induced setting.
  • The largest induced $P$-free subset of a $p$-random subset of $\mathcal{B}_{n}$ for $p\gg n^{-1}$ has size at most $(k-1+o(1))p{n\choose\lfloor n/2\rfloor}$, generalizing previous work of Balogh, Mycroft, and Treglown [4] and of Collares and Morris [10] for the case when $P$ is a chain.
All three results are asymptotically tight and give affirmative answers to general conjectures of Gerbner, Nagy, Patkós, and Vizer [18] in the case of tree posets.
@@ 1 Introduction
The celebrated Sperner’s theorem [41] in extremal set theory determines the size of the largest family of sets in $[n]$ not containing a $2$-chain $F_{1}\supset F_{2}$. Later, Erdős [16] extended Sperner’s theorem to determine the largest family not containing any $k$-chain and showed how Sperner’s lemma can be used to solve the classical Littlewood–Offord problem [28]. Afterwards, Katona and Tarján [25] initiated a systematic study of the size of the largest family in the Boolean lattice $\mathcal{B}_{n}$ that avoids a given subposet. This topic has attracted much attention and witnessed many advances in the last decades; we refer the interested reader to the nice survey paper by Griggs and Li [20] for more.
The question we are interested in this paper is one of supersaturation, that is, how many copies of a poset are we guaranteed once we are above the threshold of containing one. The simplest poset to consider the supersaturation question for are $2$-chains. Erdős and Katona conjectured that a family with $\binom{n}{\lfloor n/2\rfloor}+t$ sets in $[n]$ must contain at least $t\cdot\lfloor\frac{n+1}{2}\rfloor$ many $2$-chains. This conjecture was confirmed by Kleitman [23], who in fact showed that for every $0\leq a\leq 2^{n}$, every family in $[n]$ with size $a$ contains at least as many $2$-chains as the so-called centralized family of size $a$, i.e. a family of $a$ sets whose cardinalities are as close to $n/2+1/4$ as possible. Kleitman [23] further conjectured that the same families should also minimize the number of $k$-chains for every $k$. Five decades later, Kleitman’s result was rediscovered by Das, Gan, and Sudakov [11] and independently by Dove, Griggs, Kang, and Sereni [14]. Both papers further confirmed Kleitman’s conjecture for every $k$ and $a$ belonging to a certain range above the sum of the $k-1$ largest binomial coefficients. Subsequently, Balogh and Wagner [6] proved Kleitman’s conjecture for all $k$ and $a\leq(1-\varepsilon)2^{n}$, provided that $n$ is sufficiently large with respect to $k$ and $\varepsilon$. Finally, in a remarkable paper, Samotij [39] resolved Kleitman’s conjecture in full. There have been further generalizations of the supersaturation problem for $2$-chains to more general hosts other than the Boolean lattice such as the collection of subspaces of $\mathbb{F}_{q}^{n}$ ordered by set inclusion [32] or $\{0,1,2,\dots r\}^{n}$ [6, 32]. The latter problem gives rise to a natural generalization of Kleitman’s problem, and while weaker approximate results do hold for $r\geq 2$ as shown by Noel, Sudakov and Scott [32], the exact analogue of Kleitman’s conjecture (see  [6, 32]) fails as shown by Balogh, Petříčková, and Wagner [7].
Our main goal is to establish similar supersaturation results for more general family of posets beyond 2-chains, and in particular we do this for so-called tree posets, which are posets whose Hasse diagram is a tree. To state our results, we need a short prelude on the precise definition of extremal numbers for posets, since below this extremal number we can not guarantee any copies of our poset.
Let P, Q be two finite posets, that is, they are finite sets equipped with partial orders $<_{P}$ and $<_{Q}$. A poset homomorphism is a function $f:P\to Q$ such that $f(x)<_{Q}f(y)$ whenever $x<_{P}y$. An induced poset homomorphism is a function $f:P\to Q$ such that $f(x)<_{Q}f(y)$ if and only if $x<_{P}y$. We say that a poset $Q$ contains another poset $P$ if there is an injective poset homomorphism from $P$ to $Q$. We say that a poset $Q$ contains an induced copy of another poset $P$ if there is an injective induced poset homomorphism from $P$ to $Q$. If a poset $Q$ does not contain a copy of another poset $P$, we say that $Q$ is $P$-free. If $Q$ does not contain an induced copy of $P$, we say that $Q$ is induced $P$-free. Given a poset $P$ and an integer $n$, we define $La(n,P)$ to be the largest size of a $P$-free subfamily of $\mathcal{B}_{n}$ and $La^{*}(n,P)$ the largest size of an induced $P$-free subfamily of $\mathcal{B}_{n}$. Motivated by a number of early results in the study of $La(n,P)$ (see for example [12, 13, 19, 42]), Griggs and Lu [21] and independently Bukh [9] made the following conjecture on the form of $La(n,P)$.
@@
A similar conjecture for the induced case as well as a supersaturation version of both results were stated in Gerbner, Nagy, Patkós, and Vizer (see Conjecture 1 and Conjecture 3  [18]). An approximate version of Conjecture 1.1 was proven by Methuku and Pálvölgyi [30] who showed that for every poset $P$, there exists a constant $C_{P}$ such that $La^{*}(n,P)\leq C_{P}\binom{n}{\lfloor n/2\rfloor}$. The value of $C_{P}$ was later improved by Tomon [43]. However, this conjecture as well as the companion ones from [18] were proven to be false for $P$ the $d$-dimensional Boolean lattice with $d\geq 4$ due to a recent beautiful construction of Ellis, Ivan, and Leader [15].
Nevertheless there are still a number of families of $P$ for which Conjecture 1.1 remains true (see [20] for more), the most general family perhaps being that of tree posets as established by Bukh [9].
@@
Bukh’s result was later extended to the induced setting by Boehnlein and Jiang [8], who showed that $La^{*}(n,P)=(k-1+O(\frac{\sqrt{n\log n}}{n}))\binom{n}{\lfloor n/2\rfloor}$ for any tree poset of height $k$. Our main result is the following supersaturation extension of these results.
@@
Let $k$ be a fixed positive integer and $P$ a tree poset of height $k$. Then for any real $\varepsilon>0$ and integer $q\geq k$, there exists a real $\delta>0$ such that every family $\mathcal{F}\subseteq\mathcal{B}_{n}$ with $|\mathcal{F}|\geq(q-1+\varepsilon)\binom{n}{\lfloor n/2\rfloor}$ contains at least $\delta\cdot M^{*}(n,q,P)$ induced copies of $P$ where $M^{*}(n,q,P)$ denotes the number of induced copies of $P$ in the $q$ middle levels of $\mathcal{B}_{n}$.
This result answers [18, Conjecture 3] in a strong form for tree posets which stated their conjecture only in the special case of $q=k$. In fact, an analog of Theorem 1.3 holds if we replace $|\mathcal{F}|$ with $\mathcal{F}$’s Lubell weight, see Theorem 5.4 for the precise statement. It is worth noting that we do not know of an explicit formula for $M^{*}(n,q,P)$ in general, but this does not end up being a significant barrier to the proof (see Section 3 for more details). However, as a corollary to Theorem 1.3, we can get the following explicit result which is tight for saturated tree posets, i.e. those for which every maximal chain has the same length.
@@
Let $k$ be a fixed positive integer and $P$ a tree poset of height $k$. For any real $\varepsilon>0$, there exists a real $\delta>0$ such that every subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ with $|\mathcal{F}|\geq(k-1+\varepsilon)\binom{n}{\lfloor n/2\rfloor}$ contains at least $\delta n^{|P|-1}\binom{n}{\lfloor n/2\rfloor}$ induced copies of $P$.
A slight modification of our approach can extend Corollary 1.4 into a balanced supersaturation result, Theorem 5.6, which roughly speaking says that we can guarantee our collection of copies of $P$ to be such that no subset of $\mathcal{F}$ is contained in too many induced copies of $P$ in this collection. Balanced supersaturation while interesting on its own is usually used in combination with the container method to establish counting results. The container method originated in papers by Kleitman and Winston [26, 27] in early 1980s and was further independently developed by Balogh, Morris, and Samotij [2] and Saxton and Thomason [40], and has had a tremendous impact on Combinatorics since then. Containers have been widely used to establish counting results and random Turán-type in various settings such as for $H$-free graphs or hypergraphs for a fixed graph $H$, AP-free sets in additive combinatorics and so on. See the excellent survey by Balogh, Morris and Samotij [3] for an extensive overview on the method.
In the setting of posets, containers have been most widely used in counting subsets of $\mathcal{B}_{n}$ avoiding $2$-chains, which is equivalent to the problem of counting antichains. When the host is the Boolean lattice, this problem was solved independently by Balogh, Treglown, and Wagner [5] and by Noel, Sudakov, and Scott [32]. The second group [32] also generalized this to hosts being subspaces of $\mathbb{F}_{n}^{q}$ ordered by inclusion, and sets of divisors of a square of a square free integer. Note that the latter is equivalent to studying $\{0,1,2\}^{n}$ under the natural ordering. In the more general setting of counting antichains in $[t]^{n}$, which turns out to be connected to a Ramsey-theoretic question in ordered hypergraphs [31], there has been much recent progress, most notably by Pohoata and Zakharov [37], Park, Sarantis, and Tetali [34] and finally, by Falgrav-Ravry, Rävy, and Tomon [17].
Our main counting result gives tight bounds for the number of induced $P$-free families of $\mathcal{B}_{n}$ for all tree posets $P$, solving Conjecture 4 from [18] in this case. This extends previous work of Patkós and Treglown [36] and Gerbner, Nagy, Patkós, and Vizer [18] who obtained similar results for special subclasses of tree posets. Our result also strengthens the recent work of Balogh, Garcia, and Wigal [1] who independently obtained the same counting result but in the non-induced setting.
Theorem 1.5.
If $P$ is a tree poset of height $k$, then the number of induced $P$-free sets in $\mathcal{B}_{n}$ is at most
$2^{(k-1+o(1))\binom{n}{n/2}}.$
With a similar approach, we also obtain tight bounds for the largest size of an induced $P$-free family of a random subset of $\mathcal{B}_{n}$ for all tree posets $P$. Let $\mathcal{P}(n,p)$ be the uniformly random subset of $\mathcal{B}_{n}$, where each set survives with probability $p$ such that $pn\rightarrow\infty$. This model $\mathcal{P}(n,p)$ was first investigated by Renyi [38], after which Kohayakawa and Kreuter [24] studied the size of the largest $2$-chain-free subset of $\mathcal{P}(n,p)$. Their results were subsequently improved by Osthus [33] and Balogh, Mycroft and Treglown [4], the latter establishing optimal bounds, and independently by Collares and Morris [10] who established the analogous results for $k$-chains for all $k\geq 2$. Other results in this direction were obtained by Patkós and Treglown [36] and by Gerbner, Nagy, Patkós, and Vizer [18] for some special subfamilies of tree posets. We establish a far-reaching generalization of all these results by establishing the corresponding result on $\mathcal{P}(n,p)$ for all tree posets $P$ in the induced setting, solving Conjecture 7 of [18] in this case.
Theorem 1.6.
If $P$ is a tree poset and $pn\rightarrow\infty$, then with high probability, the largest induced $P$-free subset of $\mathcal{P}(n,p)$ has size $(k-1+o(1))p\binom{n}{n/2}$.
The hypothesis that $pn\to\infty$ is best possible for this result to hold. Indeed, as noted in [10, 4], if $p=cn^{-1}$ then a computation due to Osthus [33] shows that with high probability there exists subsets of $\mathcal{P}(n,p)$ of size at least $(k-1+e^{-c}+o(1)){n\choose n/2}$ which contains no $k$-chain and thus also no tree poset of height $k$.
Finally, it is worth highlighting that the main tool of this paper Theorem 4.15 which we believe is of independent interest, and can be viewed as a general tool similar to results in graphs of passing to subgraphs of high minimum degree. While the exact analogue of such a result is out of reach, that is, obtaining a subfamily in the poset setting of “high minimum degree” we obtain a sequence of nested subfamilies of our original family $\mathcal{F}$ having high minimum degree in the predecessor of the sequence which still allows us to greedily find many embeddings of a tree poset, similar to embedding copies of a fixed tree in a high minimum degree subgraph. We believe that this result will have further applications in the future.
The rest of this paper is organized as follows. In Section 2 we introduce some notation. In Section 3 we present a sketch of the proof of Theorem 1.3. In Section 4 after gathering some preliminary lemmas, we prove Theorem 4.15, in Section 5 we prove both the usual and balanced supersaturation results Theorem 1.3 and Theorem 5.6. In Section 6 we use hypergraph containers together with balanced supersaturation to prove Theorem 1.5 and Theorem 1.6.